find two natural numbers whose difference is 4 and sum of their squares is 8

Grade:8

1 Answers

Santosh Kumar Thakur
6 years ago
It is not possible. Because
 
Let the two numbers be A and B where A>B.
Now, A/Q
 A^2 + B^2 =8         ….....(i)
A – B = 4                …......(ii)
 
From eqn. (ii) we get
A = 4 +B
 
Now putting the value of A in eqn (i) we get
     (4+B)^2 + B^2 = 8
=> 16 + B^2 + 8B + B^2 = 8
=> 2B^2 + 8B + 8 = 0
So, B = -2        (By Quadratic Equation)
But  -2 is not a natural number.
Hence It isn.t possible.